Cartwright Brothers’ stock is currently selling for $40 a share. The stock is expected to pay a $4 dividend at the end of the year. The stock’s dividend is expected to grow at a constant rate of 7 percent a year forever. The risk-free rate (kRF) is 7 percent and the market risk premium (kM – kRF) is 5 percent. What is the stock’s beta?

a.

5.00

b.

1.00

c.

2.00

d.

3.00

e.

4.00
Answer #1
Solution:
Answer is C. 2.00
Working Notes:
Using Gordon growth model :
P0 = D1 / (Ke - g),
ke = cost of Equity = r = ??
Po=current share price = $40 per share
g= growth rate= 7%
D1= Next Dividend=$4 per share
P0 =D1/(Ke -g)
40 = 4/(Ke-0.07)
Ke =(4/40) + 0.07
Ke =0.10+ 0.07
Ke =0.17    or 17%
We can get cost of equity using CAPM
Cost of equity
= risk-free rate + risk premium x Beta
Risk free rate kRF= 7%
Market risk premium (kM – kRF)= 5%
Beta of the stock     =b
cost of equity ke = 17% calculated above
Cost of equity
= risk-free rate + risk premium x Beta
Ke = kRF + (kM – kRF) x b
17% = 7% + 5% x b
b=(17% - 7%)/5%
b=(10%)/5%
b=2
Beta of the stock = 2
Please feel free to ask if anything about above solution in comment section of the question.